# Electromagnetic Multipole Theory for Two-dimensional Photonics

### v1

@@ 2 Validation
Figure 8: (a) TE scattering by a dimer. Values of parameters: $\epsilon_{r}=25$, $\mu_{r}=1$, $a=20~{}\mu{\rm m}$, $d=60~{}\mu{\rm m}$. (b) $Q_{\rm sc}/a$ spectrum. Curves: MAS; dots: this work via COMSOL; black: full-wave; red: MD; green: ED.
In Figure 7(a) we depict a core-shell cylinder consisting of a high-index dielectric $c$-radius circular core of relative permittivity $\epsilon_{cr}$ and relative permeability $\mu_{cr}$, coated by a gyroelectric shell of elliptical cross-section of semi-major axis $a$ and semi-minor axis $b$. Light impinges along the major axis of the ellipse in TE polarization. To calculate the full-wave $Q_{\rm sc}/a$, shown by the black curve in Figure 7(b), we employ the MAS 34. The corresponding multipole decomposition via the MAS is shown by red—MD—and green—ED—curves. In addition, the black dots correspond to the full-wave $Q_{\rm sc}/a$ as computed by eq 23, while the red and green dots correspond to the multipole decomposition using the proposed theory. The MAS and the proposed theory follow each other, and the agreement is very good.
Until now, we examined structures constituted by sole cylinders, but the theory also holds for a collection of such. As an example, we assume the dimer shown in Figure 8(a). Two identical high-index dielectric cylinders of permittivity $\epsilon_{r}$ and permeability $\mu_{r}$ are placed with a center-to-center distance $d$. Light impinges along the $x$ axis, with TE polarization, as shown in Figure 8(a). The full-wave $Q_{\rm sc}/a$ of the dimer is calculated using the MAS, and it is shown by the black curve in Figure 8(b). The multipole decomposition via the MAS is shown by red—MD—and green—ED—curves. The corresponding full-wave $Q_{\rm sc}/a$ via eq 23 is shown by black dots, and the multipole decomposition via the proposed theory by colored symbols that match perfectly the MAS solution.
@@ 3 Application on Oligomers-based Highly Directional Switching Using Active Media
The developed framework is employed to study an emerging photonics application on oligomers-based highly directional switching using active media. In Figure 9(a), we employ a dimer whose elements are made of core-shell cylinders with a center-to-center distance $d_{y}=60~{}\mu{\rm m}$. Each core has radius $b=15~{}\mu{\rm m}$ and consists of a high-index dielectric material of relative permittivity $\epsilon_{cr}=25$ and relative permeability $\mu_{cr}=1$. Each shell has an outer radius $a=20~{}\mu{\rm m}$ and consists of an active magneto-optical (MO) medium which is tunable under an external magnetic flux density ${\mathbf{B}}_{0}=B_{0}{\mathbf{e}}_{z}$. To this end, we use indium antimonide (InSb) whose gyroelectric permittivity is given by eq 1 with $\epsilon_{\rm 1r}=\epsilon_{\infty}\{1-(\omega-iv)\omega_{p}^{2}/\{\omega[(\omega-iv)^{2}-\omega_{c}^{2}]\}\}$, $\epsilon_{\rm 2r}=\epsilon_{\infty}\{\omega_{c}\omega_{p}^{2}/\{\omega[(\omega-iv)^{2}-\omega_{c}^{2}]\}\}$ and $\epsilon_{\rm 3r}=\epsilon_{\infty}\{1-\omega_{p}^{2}/[\omega(\omega-iv)]\}$ 38. In these definitions, $\epsilon_{\infty}=15.6$ accounts for interband transitions, $\omega_{p}=[N_{e}e^{2}/(\epsilon_{0}\epsilon_{\infty}m^{\ast})]^{1/2}=4\pi\times 10^{12}$ ${\rm rad~{}s}^{-1}$ is the plasma angular frequency (with $N_{e}$ the electron density, $e$ the elementary charge and $m^{\ast}=0.0142m_{e}$ electron’s effective mass, where $m_{e}$ is the electron’s rest mass), $\omega_{c}=eB_{0}/m^{\ast}$ is cyclotron angular frequency, and $v=\alpha\omega_{p}$ the damping angular frequency with $\alpha$ a dimensionless parameter. Based on the adopted time dependence $\exp(i\omega t)$, the MO material is passive (lossy) for $\alpha>0$ and active (exhibits gain) for $\alpha<0$. Therefore, by assuming the parameters of InSb, a passive material, we set negative values to the parameter $\alpha$ and make it active. In Figure 9(b), we plot the full-wave $Q_{\rm sc}/a$ spectrum, as well as the multipole decomposition, when an external $B_{0}=0.1$ T is applied, and active shells are used in the dimer with $\alpha=-0.001$. Due to the external bias, a magneto-plasmonic (MP) mode 10 is induced at $f=2.0186$ THz, whose peak is shown by the orange circle. The multipole decomposition reveals that this MP mode is due to a collective contribution between the MD and EQ responses that resonate at the same frequency. The phenomenon is further enhanced by the ED response, which is almost resonant at the same frequency. Precisely at the MP mode, we plot the full-wave normalized scattering width, i.e., $\sigma/a$, as computed by COMSOL. This full-wave simulation is depicted in the inset of Figure 9(b), where a directional forward-scattering is shown.
@@
In Figure 10(a), we plot the FOM vs various values of the dimensionless parameter $\alpha$. There exists a specific value $\alpha=-0.0007$ where the FOM is maximized. In Figure 10(b), we plot the $\sigma/a$ using this specific value of $\alpha$, as well as its opposite value, i.e., $\alpha=+0.0007$. As revealed by the zoom-in, the active shell medium ($\alpha=-0.0007$) eliminates the tail at the back-scattering direction, while the passive shell medium ($\alpha=+0.0007$) does not. Such a tail-suppression to obtain directionality is also expected in lossy structures that are illuminated using complex-frequency waves, exhibiting the so-called virtual gain, as shown in recent publications 39, 40, 41.
To further enhance the scattering and yield it highly directive, we examine in Figure 11(a) a core-shell octamer having the same unit structure and values of parameters as the dimer of Figure 9(a). Using active shells ($\alpha=-0.001$), in Figures 11(b) and (c) we plot the $Q_{\rm sc}/a$ spectra and the multipole decompositions, when $B_{0}=0.1$ T and $B_{0}=0$ T. Due to the electrically large size of the octamer, higher-order multipole responses contribute to the spectra, however these are not depicted, for clarity of the graph. Now, the MP mode (depicted by the orange circle in Figure 11(b)) takes place at $f=2.0190$ THz, which is still a collective contribution of at least the MD and EQ responses. When $B_{0}=0$ T, a smaller scattering arises at the same frequency (depicted by the orange circle in Figure 11(c)).
Finally, to optimize the forward-scattering of the octamer’s MP mode, in Figure 12(a), we show that for $\alpha=-0.0001$, the FOM becomes maximum. With this value of $\alpha$, in Figure 12(b), we plot the $\sigma/a$ when the magnetic bias is at on-state. A highly directive lobe is revealed at the forward-scattering with tail elimination at the back-scattering. When the magnetic bias is at off-state, Figure 12(c) shows a significant back-scattering with a remaining forward-scattering. Nevertheless, this on-off state suggests that a core-shell oligomer with active medium shells can operate as a highly directional active switching device in contemporary photonics.
Figure 9: (a) TE scattering by a high-index core-MO shell dimer. Values of parameters: $\epsilon_{cr}=25$, $\mu_{cr}=1$, MO shell, $a=20~{}\mu{\rm m}$, $b/a=0.75$ and $d_{y}=60~{}\mu{\rm m}$. (b) $Q_{\rm sc}/a$ spectrum for $B_{0}=0.1$ T and $\alpha=-0.001$ (active). Black: full-wave; red: MD; green: ED; blue: EQ. Inset: $\sigma/a$ at $f=2.0186$ THz.
@@
Figure 10: (a) FOM vs $\alpha$ for the dimer of Figure 9(a) when $B_{0}=0.1$ T and $f=2.0186$ THz. (b) $\sigma/a$ at $\alpha=\mp 7\times 10^{-4}$, $B_{0}=0.1$ T and $f=2.0186$ THz. Blue: $\alpha=-7\times 10^{-4}$ (active); red: $\alpha=+7\times 10^{-4}$ (passive). Zoom in: tail elimination at back-scattering.
@@
Figure 11: (a) TE scattering by a high-index core-MO shell octamer. Values of parameters: $\epsilon_{cr}=25$, $\mu_{cr}=1$, MO shell, $a=20~{}\mu{\rm m}$, $b/a=0.75$ and $d_{y}=60~{}\mu{\rm m}$. (b) $Q_{\rm sc}/a$ spectrum for $B_{0}=0.1$ T and $\alpha=-0.001$ (active). Black: full-wave; red: MD; green: ED; blue: EQ. Orange circle: MP mode at $f=2.0190$ THz. (c) Same as (b) but for $B_{0}=0$ T. Orange circle: scattering at $f=2.0190$ THz.
@@
Figure 12: (a) FOM vs $\alpha$ for the octamer of Figure 11(a) when $B_{0}=0.1$ T and $f=2.0190$ THz. (b) $\sigma/a$ at $\alpha=-1\times 10^{-4}$ (active), $B_{0}=0.1$ T and $f=2.0190$ THz. (c) Same as (b) but for $B_{0}=0$ T.
@@ 4 Conclusion
In this work, we developed and validated a comprehensive EM multipole decomposition framework for 2-D structures, with a specific focus on inhomogeneous and anisotropic cylindrical scatterers. Our method leverages the expansion of scattered fields using divergenceless CVWF and employs 2-D volumetric integrals to express the unknown expansion coefficients. Our results on multipole decomposition using finite elements simulations were validated by comparing our results with analytical and numerical methods that provide inherently the multipole decomposition. Furthermore, we demonstrated the applicability of the developed framework by analyzing photonics applications on oligomers-based highly directional-scattering switching using active MO media. Our findings reveal that the collective contributions of magnetic and electric multipole responses can be harnessed to achieve highly directional forward-scattering, a promising avenue for future photonic devices.

### v2

@@ 2 Validation
In Figure 7(a) we depict a core-shell cylinder consisting of a high-index dielectric $c$-radius circular core of relative permittivity $\epsilon_{cr}$ and relative permeability $\mu_{cr}$, coated by a gyroelectric shell of elliptical cross-section of semi-major axis $a$ and semi-minor axis $b$. Light impinges along the major axis of the ellipse in TE polarization. To calculate the full-wave $Q_{\rm sc}/a$, shown by the black curve in Figure 7(b), we employ the MAS 42. The corresponding multipole decomposition via the MAS is shown by red—MD—and green—ED—curves. In addition, the black dots correspond to the full-wave $Q_{\rm sc}/a$ as computed by eq 23, while the red and green dots correspond to the multipole decomposition using the proposed theory. The MAS and the proposed theory behave similarly and the agreement is satisfactory.
Figure 8: (a) TE scattering by a dimer. Values of parameters: $\epsilon_{r}=25-i2$, $\mu_{r}=1$, $a=20~{}\mu{\rm m}$, $d=60~{}\mu{\rm m}$. (b) $Q_{\rm sc}/a$ spectrum. Curves: MAS; dots: this work via COMSOL; black: full-wave; red: MD; green: ED. (c) Same as (b) but for $Q_{\rm ext}/a$.
Until now, we examined structures constituted by sole cylinders, but the theory also holds for a collection of such. As an example, we assume the dimer shown in Figure 8(a). Two identical high-index lossy dielectric cylinders of permittivity $\epsilon_{r}$ and permeability $\mu_{r}$ are placed with a center-to-center distance $d$. Light impinges along the $x$ axis, with TE polarization, as shown in Figure 8(a). The full-wave $Q_{\rm sc}/a$ of the dimer is calculated using the MAS, and it is depicted by the black curve in Figure 8(b). The multipole decomposition via the MAS is shown by red—MD—and green—ED—curves. The corresponding full-wave $Q_{\rm sc}/a$ via eq 23 is shown by black dots, and the multipole decomposition via the proposed theory, by colored symbols that match perfectly the MAS solution. Since the cylinders are lossy, in Figure 8(c), we further calculate the full-wave and the multipole decomposition quantities for the $Q_{\rm ext}/a$. In particular, the full-wave $Q_{\rm ext}/a$ in COMSOL is computed by $Q_{\rm ext}/a=Q_{\rm sc}/a+Q_{\rm abs}/a$ via eqs 23 and 25, while the $Q_{\rm ext}/a$ multipole decomposition from the proposed theory via eq 26. As evident, all quantities are in agreement.
@@
Figure 9: (a) TE scattering by a high-index core-MO shell dimer. Values of parameters: $\epsilon_{cr}=25$, $\mu_{cr}=1$, MO shell, $a=20~{}\mu{\rm m}$, $b/a=0.75$ and $d_{y}=60~{}\mu{\rm m}$. (b) $Q_{\rm sc}/a$ spectrum for $B_{0}=0.1$ T and $\alpha=-0.001$ (active). Black: full-wave; red: MD; green: ED; blue: EQ. Inset: $\sigma/a$ at $f=2.0186$ THz. (c) $Q_{\rm sc}/a$, $Q_{\rm abs}/a$, and $Q_{\rm ext}/a$ spectra for the same values of parameters as in (b). Black: full-wave $Q_{\rm sc}/a$ as computed by eq 23; cyan curve: full-wave $Q_{\rm abs}/a$ as computed by eq 25; orange: full-wave $Q_{\rm ext}/a$ as computed by eq 26; cyan dots: $Q_{\rm ext}/a-Q_{\rm sc}/a$.
@@ 3 Application on Oligomers-based Highly Directional Switching Using Active Media
The developed framework is employed to study an emerging photonics application on oligomers-based highly directional switching using active media. In Figure 9(a), we employ a dimer whose elements are made of core-shell cylinders with a center-to-center distance $d_{y}=60~{}\mu{\rm m}$. Each core has radius $b=15~{}\mu{\rm m}$ and consists of a high-index dielectric material of relative permittivity $\epsilon_{cr}=25$ and relative permeability $\mu_{cr}=1$. Each shell has an outer radius $a=20~{}\mu{\rm m}$ and consists of an active magneto-optical (MO) medium which is tunable under an external magnetic flux density ${\mathbf{B}}_{0}=B_{0}{\mathbf{e}}_{z}$. To this end, we use indium antimonide (InSb) whose gyroelectric permittivity is given by eq 1 with $\epsilon_{\rm 1r}=\epsilon_{\infty}\{1-(\omega-iv)\omega_{p}^{2}/\{\omega[(\omega-iv)^{2}-\omega_{c}^{2}]\}\}$, $\epsilon_{\rm 2r}=\epsilon_{\infty}\{\omega_{c}\omega_{p}^{2}/\{\omega[(\omega-iv)^{2}-\omega_{c}^{2}]\}\}$ and $\epsilon_{\rm 3r}=\epsilon_{\infty}\{1-\omega_{p}^{2}/[\omega(\omega-iv)]\}$ 48. In these definitions, $\epsilon_{\infty}=15.6$ accounts for interband transitions, $\omega_{p}=[N_{e}e^{2}/(\epsilon_{0}\epsilon_{\infty}m^{\ast})]^{1/2}=4\pi\times 10^{12}$ ${\rm rad~{}s}^{-1}$ is the plasma angular frequency (with $N_{e}$ the electron density, $e$ the elementary charge and $m^{\ast}=0.0142m_{e}$ electron’s effective mass, where $m_{e}$ is the electron’s rest mass), $\omega_{c}=eB_{0}/m^{\ast}$ is the cyclotron angular frequency, and $v=\alpha\omega_{p}$ the damping angular frequency with $\alpha$ a dimensionless parameter. Based on the adopted time dependence $\exp(i\omega t)$, the MO material is passive (lossy) for $\alpha>0$ and active (exhibits gain) for $\alpha<0$. Therefore, by assuming the parameters of InSb, a passive material, we set negative values to the parameter $\alpha$ and make it active. In Figure 9(b), we plot the full-wave $Q_{\rm sc}/a$ spectrum, as well as the multipole decomposition, when an external $B_{0}=0.1$ T is applied, and active shells are used in the dimer with $\alpha=-0.001$. Due to the external bias, a directional-scattering mode 11 is induced at $f=2.0186$ THz, whose peak is shown by the orange circle. The multipole decomposition reveals that this mode is due to a collective contribution between the MD and EQ responses that resonate at the same frequency. The phenomenon is further enhanced by the ED response, which is almost resonant at the same frequency. Precisely at the peak, we plot the full-wave normalized scattering width, i.e., $\sigma/a$, as computed by COMSOL. This full-wave simulation is depicted in the inset of Figure 9(b), where a directional forward-scattering is shown. Since the structure in Figure 9(a) incorporates an active medium, we discuss the specifics of the scattering from the point of view of energy relations, by calculating the absorption and the extinction cross-section. In Figure 9(c), we plot the full-wave $Q_{\rm sc}/a$, the full-wave $Q_{\rm abs}/a$ and the full-wave $Q_{\rm ext}/a$ spectra, as computed by eqs 23, 25, and 26, respectively. The difference $Q_{\rm ext}/a-Q_{\rm sc}/a$—cyan dots in Figure 9(c)—is in perfect agreement with $Q_{\rm abs}/a$—cyan curve in Figure 9(c)—proving the correctness of the calculations. The $Q_{\rm abs}/a$ is negative, indicating emission due to the active layers. Therefore, the energy scattered by the cylinder equals the energy provided by the source via the incident plane wave, plus the energy due to the emission. In the spectral window of Figure 9(c), the emerging mode at $f=2.0186$ THz exhibits the maximum gain.
@@
In Figure 10(a), we plot the FOM vs. various values of the dimensionless parameter $\alpha$. There exists a specific value $\alpha=-0.0007$ where the FOM is maximized. In Figure 10(b), we plot the $\sigma/a$ using this specific value of $\alpha$, as well as its opposite value, i.e., $\alpha=+0.0007$. As revealed by the zoom-in, the active shell medium ($\alpha=-0.0007$) eliminates the tail at the back-scattering direction, while the passive shell medium ($\alpha=+0.0007$) does not. Such a tail-suppression to obtain directionality is also expected in lossy structures that are illuminated using complex-frequency waves, exhibiting the so-called virtual gain, as shown in recent publications 49, 50, 51.
@@
Figure 10: (a) FOM vs. $\alpha$ for the dimer of Figure 9(a) when $B_{0}=0.1$ T and $f=2.0186$ THz. (b) $\sigma/a$ at $\alpha=\mp 7\times 10^{-4}$, $B_{0}=0.1$ T and $f=2.0186$ THz. Blue: $\alpha=-7\times 10^{-4}$ (active); red: $\alpha=+7\times 10^{-4}$ (passive). Zoom in: tail elimination at back-scattering.
@@
Figure 11: (a) $Q_{\rm sc}/a$ spectrum for the same values of parameters as in Figure 10(b) using $\alpha=-7\times 10^{-4}$. Black: full-wave; red: MD; green: ED; blue: EQ. Orange dotted arrow: location of $f=2.0186$ THz. (b) Same as (a) but for the $Q_{\rm ext}/a$ spectrum.
To further explain the origin of the FOM maximization that yields a back-scattering suppression at $f=2.0186$ THz when $\alpha=-0.0007$, we plot in Figures 11(a)–(b) the full-wave $Q_{\rm sc}/a$ and $Q_{\rm ext}/a$ spectra, as well as their multipole decompositions. As evident, since the structure of the dimer operates above the subwavelength regime, higher-order multipoles contribute in the construction of the emerging forward-scattering mode. Specifically at $f=2.0186$ THz, the FOM maximization—featuring the back-scattering suppression—is primarily contributed by the interference of the ED and EQ terms, with a secondary confluence from the MD term. These multipole terms interfere destructively toward the back-scattering direction of Figure 10(b), producing the FOM maximization.
Figure 12: (a) TE scattering by a high-index core-MO shell octamer. Values of parameters: $\epsilon_{cr}=25$, $\mu_{cr}=1$, MO shell, $a=20~{}\mu{\rm m}$, $b/a=0.75$ and $d_{y}=60~{}\mu{\rm m}$. (b) $Q_{\rm sc}/a$ spectrum for $B_{0}=0.1$ T and $\alpha=-0.001$ (active). Black: full-wave; red: MD; green: ED; blue: EQ. Orange circle: Resonant mode at $f=2.0190$ THz. (c) Same as (b) but for $B_{0}=0$ T. Orange circle: scattering at $f=2.0190$ THz.
To enhance the scattering and yield it highly directive, we examine in Figure 12(a) a core-shell octamer having the same unit structure and values of parameters as the dimer of Figure 9(a). Using active shells ($\alpha=-0.001$), in Figures 12(b) and (c) we plot the $Q_{\rm sc}/a$ spectra and the multipole decompositions, when $B_{0}=0.1$ T and $B_{0}=0$ T. Due to the electrically large size of the octamer, higher-order multipole responses contribute to the spectra, however these are not depicted for graph clarity. Now, the peak (depicted by the orange circle in Figure 12(b)) takes place at $f=2.0190$ THz, which is still a collective contribution of at least the MD and EQ responses. When $B_{0}=0$ T, a smaller scattering arises at the same frequency (depicted by the orange circle in Figure 12(c)).
@@
Figure 13: (a) FOM vs. $\alpha$ for the octamer of Figure 12(a) when $B_{0}=0.1$ T and $f=2.0190$ THz. (b) $\sigma/a$ at $\alpha=-1\times 10^{-4}$ (active), $B_{0}=0.1$ T and $f=2.0190$ THz. (c) Same as (b) but for $B_{0}=0$ T.
Finally, to optimize the forward-scattering of the octamer’s resonant mode, in Figure 13(a), we show that for $\alpha=-0.0001$, the FOM becomes maximum. With this value of $\alpha$, in Figure 13(b), we plot the $\sigma/a$ when the magnetic bias is at on-state. A highly directive lobe is revealed at the forward-scattering with tail elimination at the back-scattering. When the magnetic bias is at off-state, Figure 13(c) shows a back-scattering with a remaining forward-scattering. Nevertheless, this on-off state suggests that a core-shell oligomer with active medium shells can operate as a highly directional active switching device in contemporary photonics.
@@ 4 Conclusion
In this work, we developed and validated a comprehensive EM multipole decomposition framework for 2-D structures, with a specific focus on inhomogeneous and anisotropic cylindrical scatterers. Our method leverages the expansion of scattered fields using divergenceless CVWF and employs 2-D volumetric integrals to express the unknown expansion coefficients. Our results on multipole decomposition using finite element simulations were validated by comparing our results with analytical and numerical methods that provide inherently the multipole decomposition. Furthermore, we demonstrated the applicability of the developed framework by analyzing photonic applications on oligomers-based highly directional-scattering switching using active MO media. Our findings reveal that the collective contributions of magnetic and electric multipole responses can be harnessed to achieve highly directional forward-scattering, a promising avenue for future photonic devices.